OpenStudy (anonymous):

Solve for w in P = 2w + 2l, if P = 38 and l = 12. w =___________ a0 please help me:(

OpenStudy (anonymous):

plug in p=38, and I=12 into your equation. so, 38=2w + 2(12)

OpenStudy (anonymous):

38=2w+24 Solve for W

OpenStudy (anonymous):

so its 12???????????

OpenStudy (anonymous):

not quite

OpenStudy (anonymous):

38-24=2w 14=2w w=14/2 w=?

OpenStudy (anonymous):

so its 12???

OpenStudy (anonymous):

7

OpenStudy (anonymous):

ow haha thank you so much:)